OpenStudy (anonymous):

The Greenville City Garage charges $5 for the first hour and then $2.75 for each additional hour. How long has a car been parked at this garage if the parking charge is $29.75?

OpenStudy (anonymous):

5 + 2.75(h-1) = 29.75 2.75(h-1) = 24.75 2.75h-2.75 = 24.75 2.75h = 27.5 h = 10

OpenStudy (lgbasallote):

let x be the additional hours so the equation would be 5 + 2.75x = 29.75 subtract 5 from both sides 2.75x = 24.75 divide both sides by 2.75 x = 9 however, x is just the additional hour. the car was parked an hour before that so add 1 x + 1 = 10 hours

OpenStudy (lgbasallote):

just another way to solve it